If 0.10M solution of a weak base has a pH of 8.90, find the Kb for the base.
Answer #1

KOH Kt + он то-lo Moo E O.10 - и их Now, poн = 14-рН = 1 - &• 0 = $. (о - Сон:]: fоr thе tо -1°, 2, 34 х 10-“мэх Now, K, Ск+]